solve using the substitution method. Answer in solution set.
2x+y=0
6x-3y=12

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(57967)

solve using the substitution method. Answer in solution set.
2x+y=0
6x-3y=12
-----
y = -2x
Substitute:
6x - 3(-2x) = 12
6x + 6x = 12
12x = 12
x = 1
---------
Solve for "y":
y = -2x
y = -2*1 = -2
===========================
Cheers,
Stan H.
=================

Answer by solver91311(17077)



Solve the first equation for which is to say, add to both sides of the first equation.

Once you have an expression in that is equal to , substitute it into the second equation in place of the . Solve the resulting single variable equation for .

Once you have a value for , substitute that value in place of in either equation. Solve the resulting single variable equation in for .

Express your answer as an ordered pair of the form using the values determined in the steps above.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism